OpenStudy (anonymous):

Please help! Find the surface area generated by revolving y=2sqrt(x) From x=0 to x=8 about the x-axis.

OpenStudy (anonymous):

so you need to find the arc length between 0 and 8

OpenStudy (anonymous):

this is equal to: \[\int\limits_{0}^{8} \sqrt{dx^2 + dy^2}\]

OpenStudy (anonymous):

or\[\int\limits_{}^{} ds\]

OpenStudy (anonymous):

because: \[ds^2 = dx^2 + dy^2\]

OpenStudy (anonymous):

true even for curves at very small distances

OpenStudy (anonymous):

now: \[\int\limits_{0}^{8} \sqrt{1 + (\frac{dy}{dx})^2} dx\]

OpenStudy (anonymous):

by multiplying by: \[dx/dx\]

OpenStudy (anonymous):

then squaring it once brought inside the radical

OpenStudy (anonymous):

so take the derivative of: \[2x^{1/2}\] to get: \[x^{-1/2}\]

OpenStudy (anonymous):

square: 1/x

OpenStudy (anonymous):

add one to and square root: \[\sqrt{x^{-1}+1}\]

OpenStudy (anonymous):

integrate that with respect to dx

OpenStudy (anonymous):

either I messed up or your teacher must hate you

OpenStudy (anonymous):

haha, okay i think i got it. thank you! and i mean usually the problem he gives arnt that bad but this one about just killed me

OpenStudy (anonymous):

pluging it in to wolfram alpha, you get: http://www.wolframalpha.com/input/?i=integrate+sqrt%28x%5E-1+%2B1%29

OpenStudy (anonymous):

then you have to rotate that

OpenStudy (anonymous):

I also dont know if he wants you to find the area of the base